I HAVE 2 SEPERATE SIMILAR QUESTIONS.

Question 1)Find the equation of the line which passes through the point (0,-5)and is parralel to y=8x-3
AND
Question 2) Find the equation of the line which is perpendicular to y=5x-7 and passes through the point (0, -2)


Thank You In Advance ! :)

Answer by sudhanshu_kmr(1152)

Question 1)Find the equation of the line which passes through the point (0,-5)and is parralel to y=8x-3
slope of line = 8
(y+5) = 8(x-0)
=> y = 8x -5 (equation of required line in slope intercept form )


Question 2) Find the equation of the line which is perpendicular to y=5x-7 and passes through the point (0, -2)
slope of line = 5 , slope of required line = -1/5
(y +2) = -1/5 (x-0)
=> y + 2 = -x/5
=> y = -x/5 - 2
(equation of required line in slope intercept form )